Jake can complete a job in 5 hours while Jerry can complete the same job in 3 hours.
a) How long would it take to complete the job if both Jerry and Jake worked
together?
b) How long would it take to complete the job if Jake starts the job two hours before
Jerry joins him?
I don't know how to set either question up

Answer by stanbon(75887)

Jake can complete a job in 5 hours while Jerry can complete the same job in 3 hours.
a) How long would it take to complete the job if both Jerry and Jake worked
together?
Jake rate = 1/5 job/hr
Jerry rate = 1/3 job/hr
Together rate = 1/x job/hr
------
Equation:
rate + rate = together rate
1/5 + 1/3 = 1/x
3x + 5x = 15
8x = 15
x = 15/8 = 1 7/8 hrs = 1hr+(5/8)60 min = 1hr 52.5 minutes
==============================================================
b) How long would it take to complete the job if Jake starts the job two hours before Jerry joins him?
---
Jake will complete 2(1/5) = 2/5 of the job.
-----
Solve:
(1/5)x + (1/3)x = 3/5 job
----
3x + 5x = 9
8x = 9
x = 9/8 hrs = 1 hr 7.5 minutes
=================================
Cheers,
Stan H.